1. What are buffers?

________________________________________________________________________

________________________________


2. Write equations to show the dissociation of the salts (i) ammonium chloride and (ii) sodium acetate.

I. __________________________________________________________________

__________________________________________________________________


II. __________________________________________________________________

__________________________________________________________________


3. Write equations to show the hydrolysis of the two salts named in Exercise 2.

I. __________________________________________________________________


II. ________________________________________________________________


4. Calculate the pH of the ammonium chloride solution of concentration 0.1M if the acid dissociation constant/ hydrolysis constant for the NH4+ ion is 5.8 x 10-10

1
Expert's answer
2022-05-25T06:09:02-0400

1.

Buffer is, solution usually containing an acid and a base, or a salt, that tends to maintain a constant hydrogen ion concentration. A buffer consists of a weak acid and its conjugate base or a weak base and its conjugate acid. Buffer capacity is the amount of acid or base that can be added before the pH of a buffer changes. An example of a buffer solution is bicarbonate in blood, which maintains the body's internal pH.

2.

(i) NH4Cl => NH4+ + Cl-

(ii) CH3COONa => Na+ + CH3COO-

3.

(i) NH4Cl + H2O <=> NH4OH + HCl

(ii) CH3COONa + H2O <=> CH3COOH + NaOH

4.

a)

Kdis – constant of dissociation = 5,8 * 10-10 ;

CM – concentration of molarity = 0,1 mol/l ;

CM(NH4+) – concentration of molarity of NH4+ = X mol ;

CM(NH4+)= CM * Kdiss = 0,1 * 5,8*10-10 = 5,8*10-11

b) pH of solution: It is defined as negative of logarithm of hydrogen ion concentration. Mathematically:- pH = - log[H+] .

Substituting 5,8*10-11 into the equation for pH we find that the 0,1 M NH4Cl solution has a pH of -log(5,8*10-11) = 11-log(5,8) = 11 - 0,76 = 10,24

Answer: pH= 10,24